A liquid occupies 110 cm3 at a temperature of 12C and 117.64 cm3 at a temperature of 68C.What is coefficient of volumetric expansion for the liquid?

a) 10.21x 10-4

b) 12.40 x 10 -4

c) 6.208 x 10-4

d) 5.105 x 10-4
Answer #1

Initial volume of the liquid = V1 = 110 cm3

Initial temperature of the liquid = T1 = 12 oC

Final volume of the liquid = V2 = 117.64 cm3

Final temperature of the liquid = T2 = 68 oC

Temperature difference = \Delta T

\DeltaT = T2 - T1

\DeltaT = 68 - 12

\DeltaT = 56 oC

Coefficient of volumetric expansion for the liquid = \alpha

V2 = V1(1 + \alpha \DeltaT)

117.64 = 110(1 + \alpha (56))

1.06945 = 1 + 56\alpha

\alpha = 12.40 x 10-4  oC-1

Coefficient of volumetric expansion for the liquid = 12.40 x 10-4oC-1